FEL Upgrade Project Weekly Brief
April 2-6, 2001

The FEL team spent the week focussing on upgrade activities.  A meeting was held in the FEL vault with key WBS managers to further detail the steps to be performed on the rip-out and re-installation next fall and winter.

A talk covering the presentation at the recent FOM FELIX review was given at the Monday Upgrade meeting.  The FELIX facility is embarking on a series of upgrades to improve their FEL performance and better accommodate user needs.  They have a very successful user program with a high quality of scientific output.  It is expected to be further enhanced if a third wiggler and linac module is added to the system  They were also encouraged to increase their pulse repetition rate and add an external pulse stacker for higher pulse energies.  Several of their users also expressed interest in coming to JLab for some of their work.

JLab hosted a group of senior managers from Brookhaven National Laboratory to explore possible collaborations in the areas of high current injectors and energy recovering linacs.  They are interested in exploring the limits of these technologies.  We agreed to jointly pursue DOE, NSF, and NIH funding to support initiatives in this area for application to cooling rings and light sources.  They are also interested in performing electron transport experiments and code validation on the Upgrade as that becomes operational. They may be able to provide some technical assistance in that effort.

WBS 3 (Beam Physics):
Initial performance analysis of the IR driver suggested the possibility of improved chromatic behavior. Minor changes in sextupole placement led to much better orbit control off momentum.  Modeling of, and fitting on, chromatic variations of propagated beam envelopes provided a significantly improved solution for matching telescopes and produced a procedure for generating large acceptance matches. When plied with sufficient beer, the beamline designer might allow that an energy recovery transport with acceptance well in excess of 10% may be possible.

WBS 4 (Injector):
The layout downstream of the gun to the light box was modeled to find the best solution for beam tube sizes. Two scenarios will be looked at closer.  Continued detailing work on the shield/cesium channels design.  Received the ion pumps and part of the flanges order for the gun chamber.  Arranged with the new W&M student to start implantation work again. A test electrode will be implanted in warm-up for the Ti electrode which will run in the FET system.

WBS 6 (RF):
Zone 3 - No progress due to missing parts.  The HVPS is due 8/15/01.

Zone 4 - Door bracket was modified to close and latch properly.  The Cross-Connects for zones 2 & 4 are now wired properly.  The wiring sheets were turned over to R. Vignato for documentation.  Marked-up drawings were given to M. Gonzales for ECO's.  The Test Data book is being prepared.  Except for #8 circulator out for repair, this zone is complete and waiting for a cryomodule.

Injector RF - Factory Acceptance Tests of the 100 kW klystrons are still scheduled for 6/20/01.  The 100 kW Circulators will be delivered to JLab early next week.  The only prospect, PSC, to upgrade the HVPS's for the 100 kW systems is still waffling on his offer.  I am starting to make up a specification for the HV power transformer so that we will be able to do this work.

Gun HVPS - Glassman has received our spare unit and is working on its upgrade.  The large torus is a long lead item, so it will several months before they will be finished and tested.

WBS 8 (I&C):
The "Alarms Handler" has been installed in the control room. A piece of software was added that looks at the Beam Loss Monitors (BLMs) and prompts the operator if an attempt is made to open the drive laser shutter while the machine is in Beam Mode 3, or 4 (unsafe condition). These updates will help the transition of operations from the FEL core group to "fresh" operators. (Not "inexperienced" since they have a great deal of experience with CEBAF)  The server hard drives were also upgraded.

Design activity for a VME based Beam Loss Monitor (BLM)  card/system is ramping up. Informal discussions are ongoing regarding addition features. The Integrator "front-end" electronic will remain essentially untouched but an additional integrator will be added for a "low beam loss" diagnostic. This second integrator will be controlled by a DSP, the desire is to have a loss calculated and displayed in units something like "nanoamp-seconds". The design of the new four channel VME timing/delay module continues, this will have the VME interface placed on a "daughter-board" such that new VME designs will require only the connector be placed on the PCB layout and design efforts can be focused on the functionality of the board and not on the interface. The design of the MPS Input card is proceeding as well.

James and Dan have been working the LSS box in Lab 2.  The smoke detector is wired into the LSS box.  Still need the Altera chip programmed and some additional sensor work.  The smoke detector has been installed in the hutch in the drive laser clean room and wired into LSS as well.  Several cable connectors in the Laser enclosure were re-worked for improved performance and durability. Initial planning has gone into interfacing the YAG laser in Lab 2 for the Optics group.

The vacuum system layout Rev. 0 is complete. Drawing are being prepared then they will be reviewed. The process of gathering parts for the Ion Pump Power Supplies is nearly complete. Received bids from Allied and Newark on parts that are still needed.

Bid requests have been sent out to 5 Contractors for assembly of the User power supplies for the labs. This will also help to establish a good source for other up-coming fabrication jobs. The remaining VME front panels that were sent out for machining and silkscreen are back. All spare parts for the VIDMAR in Lab 5 have been ordered.

Drawings into EECAD:  LSS Key Shutter Control Box
        General 64 Bit Digital I/O - GPVS64 Digital I/O card schematic
        Picomotor Power Supply Chassis (For final correction)
        New OBPM Amplifier

The autotransformers arrived, these are to act as a back-up in case of a failure of the 50KWatt HPA variacs. These were responsible for a number of hours of down time during the last run.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research used their experience with their RADIA magnetic analysis code, their newly upgraded,
    high memory computers and a new element generation algorithm to investigate the field flatness and
    integral flatness on the large injector dipole (GU).  They achieved success, obtaining values for both
    characteristics of less than 1 part in 10,000 over the good field region.  In addition, the noise in their
    analysis, the mathematically generated oscillations were all below these values.  At week’s end they are
    doing one more run with length adjustment and some slight mu metal edge adjustment in search of
    flatness over a wider zone and homing in on the correct effective length.  With this result, they will start
    morphing their existing drawings of the GV to the GU.  With the insight gained from the successful
    pursuit of GU field flatness, they can turn back to the frustrating "all-ends" GV magnet and try to get its
    values to the same specification.
Optical Chicane Dipole (DW)
o We are re-checking the drawing set.
o  We are buying sample strips from five rolls of silicon steel that is destined to be used as the pole tips on the
    optical chicane dipoles and the 180° dipoles.  We will measure their thickness over the very flat center of
    the sheet and select the best roll from which to extract our pieces.  (They have to be flat to tenths of
    thousandth of an inch).
o  Leigh Harwood checked the specification for the major dipole steel purchase so that the purchase
    process can now start.
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing the 180 degree dipole (GY). They laid out the leads, hoses, manifolds
    and terminal strips for the path length correctors and brought their coil drawings to the width in the
    magnetic model.
o  The magnetic model of the GY with path length correctors on showed that their was no effect on flatness
    properties.  They also ran a number of runs with decreased Purcell gap, bringing the overall field slightly
    higher to compensate for the decreased field in the gap at the path length corrector coil slot.  There was no
    effect on the field flatness and field integral flatness.  (Which meet Dave’s specification) This may not be
    the correct fix.  David Douglas is modeling the magnet as a series of segments in DIMAD.
QG (3 inch quad)
The coils manufacturing specification was checked and the coil drawings signed in anticipation of sending the coil package to Procurement next week.
QX (Trim Quad)
o We continued to magnetic design.
o We continued laying out the girders in 3D IDEAs format and started laying out the girders and stands
o We continued laying out the beam chamber in the arcs.

WBS 10 (Wiggler):
A preliminary set of magnet measurements were performed on the dispersion section and verified the first order field requirements.  The magnet must be disassembled to add holes that the vendor missed on the support stand.  Detailed magnet measurements will start once the magnet is reassembled.

Estimates were received on the magnetic modeling of the wigglers.  The estimates were very high so it was decided to figure out the strength of the end coils by direct measurements.

We decided to analyze the end coils of the Wiggler ourselves because the cost to have AES do it was much greater than simply winding conductor on the poles and deducing the proper number of turns from
measurement (Analog computer method).

WBS 11 (Optics):
Meetings were held early in the week to discuss mirror mounting concepts for the optical cavity assemblies. The group concurred with our scheme for employing 4 mirrors mounted in JLab-designed mounts, and so we're now designing them.  An internal document was distributed for comment that details the design evolution of the optical cavity. The ultraviewer (alignment) design for 4" tubing is being detailed. Procurement received three bids to build two prototype deformable mirrors; we anticipate an award early next week. A meeting was held to discuss how best to accommodate ultrashort x-ray production and delivery in the Upgrade, and still accommodate optical and electron-beam diagnostics downstream of the wiggler. The impact for optical diagnostics was that we will no longer plan on fastening our diagnostic onto the wiggler vacuum chamber, but will instead design it on a cross.

The 50% mirror in the User Lab 1 mirror cassette was realigned to make it's beam coincident with the other two. Other vacuum parts needed for the User Lab 2 mirror cassette were received. I & C, in conjunction with the electrical installation group, wired smoke detectors in the drive laser enclosure, and in User Lab 2. Other LSS work also occurred.

A training meeting was held on the optical systems with the MCC operations coordinator to explain functionality and characteristics of the cavity and optical transport system.

    FEL X-ray Research:

    1. Boyce prepared an LOP for the next run. There are two objectives: a) measure spatial distribution
        and b)investigate the anamolous x-ray peak found in the last run.
    2. In order to accomplish the above, some equipment needs to be modified and I have designed a Pb shield
        for the X-ray detector. I've scrounged two Pb bricks that have never been exposed to radiation and
        supplied the shop with them along with the design. Casey anticipates the job to be complete by
        next Wednesday.
    3. In parallel with the shop activities, we're working to modify the detector translator station that will move
        the x-ray detector across the Be window through which the diffracted x-rays emerge.
    4. Met with Steve Benson, Michelle Shinn, and Dave Douglas to discuss the X-ray requirements for the
        upgrade. We have decided to limit the capabilities to x-rays below 50 keV. Further, the quad triplet can
        now be moved closer to the wiggler and the x-ray chamber will be between that quad triplet and the next
        quad triplet. I am working on a design of the chamber.